# Quantum dot thermal machines – a guide to engineering

### v1

@@ VI Discussion
In the limit of electron transport governed by a single narrow energy level, which is predicted to optimise performance Mahan and Sofo (1996); Whitney (2014), the microscopic dynamics of the QD are entirely incorporated into the four electron exchange rates between it and the electrodes (see Fig.4a). Instead of considering the four rates as independent, we perform an effective change of phase-space coordinates and look into their overall magnitude, $\Gamma$, and three asymmetry coefficients: the asymmetry between the probabilities of adding an electron to the QD vs. its removal, described by the entropy difference between the two charge states involved in electron transport, $\Delta S$; the asymmetry between the tunnel coupling strengths of the QD to the two electrodes, $\gamma$; and the asymmetry between positive and negative directions of electron transport, $\alpha$, associated with detailed balance breaking.
@@
What do these results mean for optimising a heat engine based on a realistic SET? Increasing the entropy difference between the charge states, $\Delta S$ has the most profound effect on output power, even when the increase in conductance is corrected for, without a significant change in efficiency or increase in noise. This agrees with high thermoelectric performance of molecules with high degrees symmetry, and thus spatial degeneracy, such as fullerenesKim et al. (2014).
@@
In experimental settings, SETs typically have high tunnel coupling asymmetry Pyurbeeva et al. (2023), as small asymmetry in device geometry leads to exponentially great asymmetry in tunnel coupling. This is an advantage, as high values of $\gamma$ decrease noise (see Fig.6a(i)). If detailed balance is broken, due to the symmetry between $\gamma$ and $\alpha$ in Eq.20, a large $\gamma$ will make the effect of even small degrees of detailed balance breaking on the power output of the heat engine significant. Figure 7 outlines the realistic advantages that can be garnered in an SET heat engine. It shows the power-efficiency and power-noise diagrams of an SET heat engine with high tunnel coupling asymmetry ($\gamma_{L}/\gamma_{R}\approx 150$) three different degeneracy values (0, 2, and 4) and an additional small degree of detailed balance breaking (up to $\approx 30\%$).
@@
We have presented a systematic study of the effect of internal dynamics of a quantum dot on thermal machine performance for the simplest possible configuration, a single quantum dot coupled to two baths, and linear response – the simplest regime from the theoretical perspective. However, there is no reason to believe that the effect of internal microscopic dynamics should be limited to these simple cases. Optimising microscopic dynamics has been a long-standing problem in applied thermoelectrics, but has been largely (with rare exceptions Allahverdyan et al. (2004, 2010)) overlooked in the field of quantum thermodynamics. Related studies include the role of degenerate excited state in an Otto engine Allahverdyan et al. (2004, 2010), or, more recently, employing catalytic states to gain advantage in heat engine and refrigerator performance Bis (2024); Łobejko et al. (2025); Fu et al. (2025).
@@
We believe that introducing quantum dots with non-trivial microscopic dynamics to more complex systems – those with time-dependent driving, multiple quantum dots, non-thermal resources Monsel et al. (2025), etc, is topic deserving greater attention, especially as such devices approach practical applicability.
Acknowledgements
@@ Appendix C Detailed balance breaking, $G$ and $L$
$P_{1}=\\ \frac{\gamma_{L}e^{S_{1}}\alpha_{+}f(\varepsilon+\alpha T)+\gamma_{R}e^{S_{1}}\alpha_{-}f(\varepsilon-\alpha T)}{\gamma_{L}e^{S_{1}}\alpha_{+}f(\varepsilon+\alpha T)+\gamma_{L}e^{S_{0}}\alpha_{-}(1-f(\varepsilon+\alpha T))+\gamma_{R}e^{S_{1}}\alpha_{-}f(\varepsilon-\alpha T)+\gamma_{R}e^{S_{0}}\alpha_{+}(1-f(\varepsilon-\alpha T))}\\[6.0pt] =\frac{e^{S_{1}}\left(\gamma_{L}\alpha_{+}f(\varepsilon+\alpha T)+\gamma_{R}\alpha_{-}f(\varepsilon-\alpha T)\right)}{\gamma_{L}f(\varepsilon+\alpha T)\left(\alpha_{+}e^{S_{1}}+\alpha_{+}e^{-\alpha}e^{\frac{\varepsilon+\alpha T}{T}}\right)+\gamma_{R}f(\varepsilon-\alpha T)\left(\alpha_{-}e^{S_{1}}+\alpha_{-}e^{\alpha}e^{\frac{\varepsilon+-\alpha T}{T}}\right)}=\\[6.0pt] =\frac{e^{S_{1}}\left(\gamma_{L}\alpha_{+}f(\varepsilon+\alpha T)+\gamma_{R}\alpha_{-}f(\varepsilon-\alpha T)\right)}{\gamma_{L}\alpha_{+}f(\varepsilon+\alpha T)\left(e^{S_{1}}+e^{S_{0}}e^{\frac{\varepsilon}{T}}\right)+\gamma_{R}\alpha_{-}f(\varepsilon-\alpha T)\left(e^{S_{1}}+e^{S_{0}}e^{\frac{\varepsilon}{T}}\right)}=\\[6.0pt] =\frac{e^{S_{1}}}{e^{S_{1}}+e^{S_{0}}e^{\frac{\varepsilon}{T}}}=\frac{1}{1+e^{\frac{\varepsilon-T\Delta S}{T}}}=f(\varepsilon-T\Delta S)$ (50)

### v2

@@ VI Discussion
In the limit of electron transport governed by a single narrow energy band, which is predicted to optimise performance Mahan and Sofo (1996); Whitney (2014), the microscopic dynamics of the QD are entirely incorporated into the four electron exchange rates between it and the electrodes (see Fig.4a). Instead of considering the four rates as independent, we perform an effective change of phase-space coordinates and look into their overall magnitude, $\Gamma$, and three asymmetry coefficients: the asymmetry between the probabilities of adding an electron to the QD vs. its removal, described by the entropy difference between the two charge states involved in electron transport, $\Delta S$; the asymmetry between the tunnel coupling strengths of the QD to the two electrodes, $\gamma$; and the asymmetry between positive and negative directions of electron transport, $\alpha$, associated with detailed balance breaking.
@@
What do these results mean for optimising a heat engine based on a realistic SET, and how can they be realised?
@@
Increasing the entropy difference between the charge states, $\Delta S$, has the most profound effect on output power, even when the increase in conductance is corrected for, without a significant change in efficiency or increase in noise. Experimentally, this can be achieved by selecting quantum dots with additional spatial degeneracy – a dimer molecule has a $\Delta S=\ln 4$ for one of the charge transitions Pyurbeeva et al. (2023) (two-fold site degeneracy in addition to spin degeneracy). Other molecules with high degrees of symmetry, and thus spatial degeneracy, such as fullerenesKim et al. (2014), can have even higher values of $\Delta S$, which agrees with their high thermoelectric performance Kim et al. (2014).
In experimental settings, SETs typically have high tunnel coupling asymmetry Pyurbeeva et al. (2023), as small asymmetry in device geometry leads to exponentially great asymmetry in tunnel coupling. This is an advantage, as high values of $\gamma$ decrease noise (see Fig.6a(i)). In solid-state devices tunnel couplings can be controlled using barrier gates, and a high asymmetry can be set by the experimentalist.
Possible avenues for realising detailed balance violation in a nanodevices include spintronics Katcko et al. (2019); Cho (2022), the inclusion of interference, such as Aharonov-Bohm-type rings Haack and Giazotto (2019); Chen et al. (2024), where the asymmetry parameter $\alpha$ could be controlled by applying a magnetic field, or, potentially, CISS Naaman et al. (2019); Bloom et al. (2024). If detailed balance breaking is present, due to the symmetry between $\gamma$ and $\alpha$ in Eq.20, a large $\gamma$ will make the effect of even small degrees of detailed balance breaking on the power output of the heat engine significant.
@@
Figure 7 outlines the realistic advantages that can be garnered in an SET heat engine. It shows the power-efficiency and power-noise diagrams of an SET heat engine with high tunnel coupling asymmetry ($\gamma_{L}/\gamma_{R}\approx 150$) three different degeneracy values (0, 2, and 4) and an additional small degree of detailed balance breaking (up to $\approx 30\%$).
We have presented a systematic study of the effect of internal dynamics of a quantum dot on thermal machine performance for the simplest possible configuration, a single quantum dot coupled to two baths. We then identified minimum guidelines for comparing and optimising SET thermal machines with unknown internal dynamics. This study was carried out under the assumptions: linear response, weak coupling limit, and a narrow transmission energy band, the simplest regime from the theoretical perspective.
Nevertheless the derived form of $G(\varepsilon)$ and $L(\varepsilon)$ has been experimentally shown Pyurbeeva et al. (2021); Volosheniuk et al. (2025) to hold well in realistic molecular devices with non-trivial structures, even beyond the original domain of the assumptions. For instance, when $\Delta T/T\sim 1$, or when energy the level splitting becomes significant. This success stems from the generality of the effective fitting of $G(\varepsilon)$ and $L(\varepsilon)$ with Eq.17, which hold even when the fitting parameters lose their physical interpretationVolosheniuk et al. (2025).
@@
The main aim of the work was to highlight the potential benefit of engineering the internal structure of a quantum dot for optimising device performance. While our results are limited to a simple case there is no reason to believe that the effects of internal microscopic dynamics should be limited to such cases. Optimising microscopic dynamics has been a long-standing problem in applied thermoelectrics, but has been largely (with rare exceptions Allahverdyan et al. (2004, 2010)) overlooked in the field of quantum thermodynamics. Related studies include the role of degenerate excited state in an Otto engine Allahverdyan et al. (2004, 2010), or, more recently, employing catalytic states to gain advantage in heat engine and refrigerator performance Bis (2024); Łobejko et al. (2025); Fu et al. (2025).
@@
We believe that introducing quantum dots with non-trivial microscopic dynamics will lead to enhanced performance. Such complex systems, with time-dependent driving, multiple quantum dots, non-thermal resources Monsel et al. (2025), will be beneficial. These devices go beyond our simple initial assumptions, such as for strong coupling. This topic deserving greater attention, especially as such devices approach practical applicability.
Acknowledgements.
@@ Appendix C Detailed balance breaking, $G$ and $L$
$P_{1}=\frac{\gamma_{L}e^{S_{1}}\alpha_{+}f(\varepsilon+\alpha T)+\gamma_{R}e^{S_{1}}\alpha_{-}f(\varepsilon-\alpha T)}{\gamma_{L}e^{S_{1}}\alpha_{+}f(\varepsilon+\alpha T)+\gamma_{L}e^{S_{0}}\alpha_{-}(1-f(\varepsilon+\alpha T))+\gamma_{R}e^{S_{1}}\alpha_{-}f(\varepsilon-\alpha T)+\gamma_{R}e^{S_{0}}\alpha_{+}(1-f(\varepsilon-\alpha T))}\\ =\frac{e^{S_{1}}\left(\gamma_{L}\alpha_{+}f(\varepsilon+\alpha T)+\gamma_{R}\alpha_{-}f(\varepsilon-\alpha T)\right)}{\gamma_{L}f(\varepsilon+\alpha T)\left(\alpha_{+}e^{S_{1}}+\alpha_{+}e^{-\alpha}e^{\frac{\varepsilon+\alpha T}{T}}\right)+\gamma_{R}f(\varepsilon-\alpha T)\left(\alpha_{-}e^{S_{1}}+\alpha_{-}e^{\alpha}e^{\frac{\varepsilon+-\alpha T}{T}}\right)}=\\ =\frac{e^{S_{1}}\left(\gamma_{L}\alpha_{+}f(\varepsilon+\alpha T)+\gamma_{R}\alpha_{-}f(\varepsilon-\alpha T)\right)}{\gamma_{L}\alpha_{+}f(\varepsilon+\alpha T)\left(e^{S_{1}}+e^{S_{0}}e^{\frac{\varepsilon}{T}}\right)+\gamma_{R}\alpha_{-}f(\varepsilon-\alpha T)\left(e^{S_{1}}+e^{S_{0}}e^{\frac{\varepsilon}{T}}\right)}=\\ =\frac{e^{S_{1}}}{e^{S_{1}}+e^{S_{0}}e^{\frac{\varepsilon}{T}}}=\frac{1}{1+e^{\frac{\varepsilon-T\Delta S}{T}}}=f(\varepsilon-T\Delta S)$ (52)